Consider the following precipitation reaction:

2 K3PO4 (aq) + 3 Co(NO3)2 (aq) ? Co3(PO4)2 (s) + 6 KNO3 (aq)

What volume of 0.222 M K3PO4 (aq) in milliliters is needed to react with 36.75 mL of 0.250 M Co(NO3)2 (aq)?

NG 5

6. Consider the following precipitation reaction:

5 Fe2+(aq) + MnO4-(aq) + 8 H+(aq) ? 5 Fe3+(aq) + Mn2+(aq) + 4 H2O(l)

An iron sample weighing 0.214 g is converted into Fe2+(aq) and requires 31.57 mL of MnO4-(aq) according to the equation above. What is the Molarity of the MnO4-(aq) solution?
Answer #1

1)

Balanced chemical equation is:

2 K3PO4 (aq) + 3 Co(NO3)2 (aq) ? Co3(PO4)2 (s) + 6 KNO3 (aq)

Here:

M(Co(NO3)2)=0.25 M

M(K3PO4)=0.222 M

V(Co(NO3)2)=36.75 mL

According to balanced reaction:

2*number of mol of Co(NO3)2 =3*number of mol of K3PO4

2*M(Co(NO3)2)*V(Co(NO3)2) =3*M(K3PO4)*V(K3PO4)

2*0.25 M *36.75 mL = 3*0.222M *V(K3PO4)

V(K3PO4) = 27.6 mL

Answer: 27.6 mL

Only 1 question at a time please